THE PAIRED t TEST Data on resting and post-exercise pulse rates collected for 8 individuals between 19 and 22 years of age. We wish to know if there is a difference in pre-exercise and post-exer cise pulse rates. Individual Number Resting Post-exercise 108 60 70 54 54 72 101 96 136 90 78 108 102 92 118 176 4
Write out all work clearly thank you
Answer #1

From the given data, values for d = Pre exercise - Post exercise are obtained as follows:

- 28, - 30, - 8, - 54, - 48,- 20, - 17, - 80

From the d values, the following statistics are calculated:

n = 8

ar{d} = - 285/8 = - 35.625

sd = 23.6156

SE = sd/sqrt{n}

=23.6156/V8 = 8.3494

Test statistic is:

t = - 35.625/8.3494 = - 4.2668

Take alpha = 0.05

ndf = 8 - 1 = 7

From Table, critical values of t = pm 2.3646

Since calculated value of t = - 4.2668 isless than critical value of t = - 2.3646, the difference is significant. Reject null hypothesis.

Conclusion:
The data support the claim that there is a difference in pre - exercise and post - exercise pulse rates.